OpenStudy (anonymous):

Joe has 10 rolls of coins. Each roll holds 12 coins. If 5 of the rolls are full and 5 of the rolls are half full, how many coins does Joe have?

OpenStudy (anonymous):

can you start this?

OpenStudy (anonymous):

how many in the five full rolls?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

how many in one half full roll?

OpenStudy (anonymous):

the answer is90

OpenStudy (anonymous):

yes, the answer is \[5\times 12+5\times 6=60+30=90\]